OpenStudy (anonymous):

Taylor is making a mini apple pie with a radius of 3.14 in for her math class. What is the area of her mini pie?

OpenStudy (anonymous):

since the area of a circle is pi r ^2 here r =3.14 so, = pi *3.14^2 =3.142*9.8596 =30.98